OpenStudy (anonymous):

What cosine function represents an amplitude of 2, a period of 2π, a horizontal shift of π, and a vertical shift of −1? I've narrowed it down to two but I'm not sure how to solve the middle f(x) = 2 cos (x − π) − 1 f(x) = 2 cos πx − 1

Directrix (directrix):

This one has period 2 because f(x) = 2 cos πx − 1 2 π/ π = 2

OpenStudy (anonymous):

@Directrix The first 2 is the period right?

Directrix (directrix):

y = 2 cos πx − 1 The 2 in front of the cosine is the amplitude.

OpenStudy (anonymous):

@Directrix Right, and the last number is the vertical shift, how do I solve the rest?

Directrix (directrix):

The attached visual may help. http://www.purplemath.com/modules/grphtrig.htm

OpenStudy (anonymous):

@Directrix Thanks, I plugged it in, so it's f(x) = 2 cos (x − π) − 1 right?

Directrix (directrix):

Phase shift is another name for horizontal shift.

Directrix (directrix):

Look at this thread. It appears to be a discussion of your problem. http://openstudy.com/updates/53eba266e4b01789aba451ad @mrsteezyyy

Directrix (directrix):

f(x) = 2 cos (x − π) − 1 --> From @SithsAndGiggles out of options f(x) = −1 cos πx + 2 f(x) = −1 cos (x − π) + 2 f(x) = 2 cos (x − π) − 1 f(x) = 2 cos πx − 1